OpenStudy (anonymous):

Find the largest value of A such that the function f(t)=(t^3)+(2t^2)-(5t)+7 is decreasing for all t in the interval (0,A).

OpenStudy (anonymous):

I have no idea how to do this

OpenStudy (turingtest):

do you know how to find where a function is increasing or decreasing?

OpenStudy (anonymous):

ive honestly been kind of lost the past few days in my math class.

OpenStudy (anonymous):

a nice explanation would be well received :)

OpenStudy (anonymous):

All right . now a function is said to be increasing if in a range it shows positive slope and decreasing if negative. slope at any point is just the derivative of the function here df(t)/dt The funtion changes from decreasing to increasing once the slope at any point becomes zero (unless t is point of inflection where it donot change from -ve to positive or +ve to negative) so inorder to find the range (0,A) we need to find the value of A where the f'(A) = 0 this means 3t^2+4t+-5=0(substitute A for t) this leave us only one positive value of A = 1/3 * ((19)^1*2 -2)

OpenStudy (anonymous):

this leave us only one positive value of A = 1/3 * ((19)^1*2 -2) explain how you got those numbers please

OpenStudy (anonymous):

sorry it is sqrt(19) just solve the quadratic

OpenStudy (anonymous):

\[f(t)=t^3+2t^2-5t+7\] is decreasing where \[f'(t)=3t^2+4t-5<0\] so this is what we have to solve

OpenStudy (anonymous):

quadratic formula gives the zeros at \[\frac{-2\pm\sqrt{19}}{3}\] and since this is a parabola that faces up, it will be negative between the zeros and positive outside them, so it is negative over the entire interval \[(\frac{-2\sqrt{19}}{3},\frac{-2+\sqrt{19}}{3})\]

OpenStudy (anonymous):

therefore the answer to your question is \[A=\frac{-2+\sqrt{19}}{3}\]

OpenStudy (anonymous):

im not getting that answer when i use the quadratic equation

OpenStudy (anonymous):

oh ... plase try again and verify steps I got the same agian. I hope you understood the way .Answer is not so important if you know how to

OpenStudy (anonymous):

i know how to use the quadratic formula but im not getting the same numbers as you all.

OpenStudy (anonymous):

show work?

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[x=\frac{-4\pm\sqrt{4^2-4\times 3\times (-5)}}{2\times 3}\] \[x=\frac{-4\pm\sqrt{76}}{6}\] \[x=\frac{-4\pm\sqrt{19}}{6}\] \[x=\frac{-2\pm\sqrt{19}}{3}\]

OpenStudy (anonymous):

typo there, should have been \[x=\frac{-4\pm2\sqrt{19}}{6}\]

OpenStudy (anonymous):

part you might have missed was \[\sqrt{76}=\sqrt{4\times 19}=2\sqrt{19}\]

OpenStudy (anonymous):

why does the 6 in the denominator become a 3????